y=5(x+2)^2+7
for the parabola find:
the vertex,
focus,
directrix line (line y=)
length of the latux rectum

Answer by Mathtut(3670)

vertex is at (-2,7)
:
length of focus is 1/4(5)=1/20 so focus is at (-2,141/20) or(-2,7.05)
:
directrix is at y=139/20 or y=6.95
:
length of latux rectum is 4(1/20)=4/20=1/5